10 mL of a 100 ppm glyphosate solution was added to 300 g of dry topsoil in a ceramic pot. 1.00 L of water was run through the pot, and filter paper and collected below. If the measurement of the solution returned a reading of 1.256 absorbance from UV-Vis spectrophotometry, then how much roundup is still in the soil? What is the weight percent roundup to soil?
1
Expert's answer
2019-04-18T05:05:43-0400

с = (4/n) x (A/0.00606) = (4/10) x (1.256/0.00606) = 82.9 ppm

Still in soil 100 – 82.9 = 17 ppm

17 ppm / 300 g x 100% = 17x10-6 / 300 x 100% = 5.7x10-6 %